what is the difference between 126 1/4 and 78 2/3?
Answer by stanbon(75887)

what is the difference between 126 1/4 and 78 2/3?
----------------------
= (4*126+1)/4 -(3*78+2)/3
---
= 505/4 - 236/3
= (3*505 - 4*236)/12
= 571/12
---
Cheers,
Stan H.